OpenStudy (anonymous):

HELP PLEASE PLEASE PLEASE!!!! write the standard form of the equation of the parabola whose directrix is y=-4 and whose focus is at (2,2)

OpenStudy (anonymous):

\[\sqrt{(x_0-2)^2+(y_0-2)^2}=\left| y_0+4 \right|\] \[(x_0-2)^2+(y_0-2)^2=(y_0+4)^2\] simplify \[x_0-4x_0-8=12y_0\] ->\[x^2/12-1/3x-2/3=y\]

OpenStudy (anonymous):

if you need me to explain how to do this, tell me

OpenStudy (anonymous):

( x - 2)² = 12 ( y +1)